Let X ⊆ N be an infinite set of natural numbers. Let f :N → X be defined

  by

f(n+1) = min(X−{f(1),f(2),...,f(n)}) forall n∈N.

f(1) = minX,

(i) Justify the definition of f. (i.e Show that f is well defined.)

(ii) Prove that f is a strictly increasing bijection.


1
Expert's answer
2021-05-07T09:07:09-0400

(i) "f:N\rightarrow X"


"f(n+1)=\text{ min }(X-[f(1),f(2),....,f(n)]) \forall n\in N"


"f(1)=\text{ min} (X)" , As n belongs to the set of natural number So "f(n+1)" can only be get values for "n\ge 1" , Thus the given function "f:N\rightarrow X" is well defined.


(ii) "f(n+1)=\text{ min }(X-[f(1),f(2),...,f(n)] \forall n\in N"


As we get the minimum value of the function at n=0, and function gives the minimum at x=1, So the given function is strictly increasing.


Also for every different values of n we get different values means f is bijective. Hence Given function is strictly incresing bijection.